A recipe calls for 9 5
9
cups of flour. If Diane is using only half the recipe, how much flour should she use? (Enter your answer as a proper fraction or mixed number, not as an improper fraction.)

Answer by checkley77(12844)

I'll assume you mean 9 5/9.
9 5/9=(81+5)/9=86/9
(86/9)/2
86/9*1/2
86/18
4 7/9 ans.